# Fast Inference of Removal-Based Node Influence

### v2

@@ 4.3. Combined Derivation and Heuristics
The way we derive Equation 16 closely corresponds to the three kinds of node influence in Figure 3. Approximation of term $T_{1}$ entails node embeddings and gradients, which corresponds to the disappearance of $v_{r}$’s embeddings. Approximation of term $T_{2}$ contributes to the structural influence $\delta Topo_{r}$, which corresponds to the change of aggregation terms. Approximation of term $T_{3}$ contributes to extending the formula from the last GNN layer to former layers, which enlarges the neighborhood size as it increases GNN layers and spreads out the influence.
Nonetheless, we cannot compute the approximation for all nodes simultaneously. We change $\boldsymbol{f_{r}}=\sum_{i=1,i\neq r}^{N}\boldsymbol{h_{i}^{(L)}}$ to the sum of all nodes’ predictions including $v_{r}$. In this way, all the nodes can share the same $f_{r}$. This will change the gradient $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(i)}}}$, but this change might be similar to all nodes due to the following lemma.
Lemma 0 ().
@@
As shown, the increased gradient could be highly relevant to the model parameters, so it might be similar among different nodes. Though it theoretically does not work on nonlinear GNNs, our experiments have shown a satisfactory performance. Now, we can approximate the influence score for all nodes simultaneously after only one forward computation and one backpropagation.
Approximation in the link prediction task is similar. We replace $\boldsymbol{f_{r}}$ in Equation 16 which is the sum of node predictions with the sum of edge predictions.
@@ 5.2. Experiment Settings
Datasets. We conduct extensive experiments on six datasets. They include four widely used benchmark citation datasets (Cora, CiteSeer, and PubMed (Sen et al., 2008), and ogbn-arxiv (Hu et al., 2020)) and two Twitter datasets (P50 and P_20_50 (Xiao et al., 2020)). The four citation networks are homogeneous undirected graphs. A node is a paper, and an undirected edge represents a citation. The original task is to predict the research field of each paper. We add a link prediction task, and we use the dot product of two nodes’ representations plus a sigmoid function to do the prediction. The two Twitter datasets are heterogeneous directed graphs. Nodes are users and directed edges represent one of five Twitter interactions or their counterparts (e.g., be followed): follow, retweet, like, reply, and mention. It has both node classification (predicting the political leaning) and link prediction, so we use their original tasks. Table 3 lists the dataset statistics. An issue is that the trained GNN model is biased to the training-set nodes/edges. To fairly evaluate node influence, we run each experiment 5 times and cycle the data split of nodes/edges by 20% per time, giving every node/edge an equal chance to show up in training, validation, or test sets. We take the mean of the 5 results.
GNN Models. We select six representative GNN models. On the four citation datasets, we use GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2017), and GCNII (Chen et al., 2020). As the ogbn-arxiv dataset is a heated OGB benchmark, we use the SOTA model at the time we started this project, DrGAT (Zou et al., 2019), to replace GAT in the node classification task. DrGAT is an improved variant of GAT, which is equipped with a dimensional reweighting mechanism. Since the two Twitter datasets are heterogeneous, the above models can no longer be directly applied, so we use TIMME model which is proposed in the same paper as the Twitter datasets (Xiao et al., 2020). TIMME tackles three challenges on the Twitter datasets: sparse feature, sparse label, and heterogeneity. We use the hyper-parameters for DrGAT and TIMME models in their GitHub repository, since their hyper-parameters have been carefully selected. We tune the hyper-parameters for GCN, GraphSAGE, GAT, and GCNII. We also tune the hyper-parameters of each approximation method for each dataset and model.
@@ 5.3. Performance Comparison
Table 4 shows the results on the four citation datasets, and Table 5 shows the results on the two Twitter datasets. NORA outperforms the baseline methods in most cases, demonstrating its effectiveness. Among the baseline methods, the “node mask” method performs the best. It is more useful in its original design, which is to analyze the influence of a few nodes or edges. When we need to model the influence of all nodes, different nodes/edges are dominantly influenced by different nodes, so it is more difficult to optimize. The “prediction” method is greatly limited by label usage. It requires additional ground truth node influence to train, but since we only have a very small training set of 7% nodes, its potential is limited. We analyze the results of enlarging the training set in the appendix. If we increase the label usage, its performance could be better, but NORA still significantly outperforms it.
Besides, we evaluate two variations of NORA as ablation studies. If we only consider the first term $T_{1}$ and the third term $T_{3}$ in the approximation (represented as NORA-$T_{1}$ in the tables), we will only consider the task-specific influence of the embeddings but ignore the structural influence, which equals setting $k_{3}^{\prime}$ to zero in Equation 16. If we only consider the second term $T_{2}$ and the third term $T_{3}$ (represented as NORA-$T_{2}$), we will only consider the structural influence, which equals to setting $k_{3}^{\prime}$ to infinity in Equation 16. NORA outperforms both variants, demonstrating the benefit of ensembling the task-specific influence and structural influence. In some situations, only considering one of them could already have a good performance that is close to NORA. It indicates that sometimes the influence is dominated by the task-specific influence or the structural influence. Ensembling both of them provides us with the opportunity to balance between them using the hyper-parameter $k_{3}^{\prime}$.

### v3

@@ 4.3. Combined Derivation and Heuristics
An intuition is that the way we derive Equation 16 corresponds to the three kinds of node influence in Figure 3. Approximation of $T_{1}$ entails node embeddings and gradients, which corresponds to the disappearance of $v_{r}$’s embeddings. Approximation of $T_{2}$ contributes to the structural influence $\delta Topo_{r}$, corresponding to the change of aggregation terms. Approximation of $T_{3}$ contributes to extending the formula from the last layer to former layers, which corresponds to the influence spreading out, since the neighborhood size grows as the GNN layers increase.
Nonetheless, we cannot compute the approximation for all nodes simultaneously. We change $\boldsymbol{f_{r}}=\sum_{i=1,i\neq r}^{N}\boldsymbol{h_{i}^{(L)}}$ to $\boldsymbol{f}=\sum_{i=1}^{N}\boldsymbol{h_{i}^{(L)}}$, so that all the nodes share the same $f$. This will change $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(i)}}}$, but the change might be similar to all nodes due to the following lemma.
Lemma 0.
@@
The gradient change caused by adding $\boldsymbol{h_{r}^{(L)}}$ to $\boldsymbol{f_{r}}$ might highly rely on the model parameters, so it might be similar among all nodes and does not impact the influence comparison. Though it theoretically does not work for nonlinear GNNs, it could be a useful intuition. Now, we can approximate the influence score for all nodes simultaneously. After one backpropagation, we get $\boldsymbol{\frac{\partial f}{\partial h_{r}^{(i)}}}$ for every node $v_{r}\in V$. Then, we can compute the approximation based on Equation 16. These calculations are tensor calculations, which only take a few seconds on GPU.
The approximation of node influence on edges in the link prediction task is similar. We replace the sum of predicted node probabilities, $\boldsymbol{f}$, with the sum of link predictions.
@@ 5.2. Experiment Settings
Datasets. We use six representative datasets. They include four widely used citation datasets (Cora, CiteSeer, and PubMed (Sen et al., 2008), and ogbn-arxiv (Hu et al., 2020)) and two Twitter datasets (P50 and P_20_50 (Xiao et al., 2020)). The four citation networks are homogeneous undirected graphs. One node is a paper, and an edge represents citation. The original task is to predict the research field of each paper (node classification). We add a link prediction task, and we use the dot product of two nodes’ representations plus a sigmoid function to do the prediction. The two Twitter datasets are heterogeneous directed graphs. One node is a user, and an edge is one of five Twitter interactions or their counterparts (e.g., be followed): follow, retweet, like, reply, and mention. It has both node classification (predicting user’s political leaning) and link prediction, so we use their original tasks. Table 3 lists the dataset statistics. An issue is that the trained GNN model is biased to the training-set nodes/edges. To fairly evaluate node influence, we run each experiment 5 times and cycle the data split of nodes/edges by 20% per time, giving every node/edge an equal chance to show up in training, validation, or test sets. We take the mean of the 5 results.
GNN Models. We use six representative GNN models. On the four citation datasets, we use GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2017), and GCNII (Chen et al., 2020). As the ogbn-arxiv dataset is a heated OGB benchmark, we use the SOTA model at the time we started this project, DrGAT (Zou et al., 2019), to replace GAT in the node classification task. DrGAT is an improved variant of GAT, which has an additional dimensional reweighting mechanism. Since the two Twitter datasets are heterogeneous, the above models can no longer be directly used, so we use TIMME model which is proposed in the same paper as the Twitter datasets (Xiao et al., 2020). TIMME tackles three challenges on the Twitter datasets: sparse feature, sparse label, and heterogeneity. We use the original hyper-parameters for DrGAT and TIMME, since they have been carefully tuned. We tune the hyper-parameters for GCN, GraphSAGE, GAT, and GCNII. We also tune the hyper-parameters of each approximation method for each dataset and model.
@@ 5.3. Performance Comparison
Table 4 shows the results on the four citation datasets, and Table 5 shows the results on the two Twitter datasets. NORA outperforms the baseline methods in most cases. Among the baseline methods, the “node mask” method performs the best. It is more useful in its original design, which is to analyze the influence on a few nodes or edges. As we want to capture the influence of a node on the whole graph, different nodes/edges are dominantly influenced by different nodes, so it is more difficult to optimize the mask. The “prediction” method is greatly limited by label usage. It requires additional ground truth to train, but since we only have a very small training set containing 7% nodes, its potential is limited. In the appendix, we find that increasing the label usage leads to better performance, but it is still worse than NORA using 10% labels.
Besides, we evaluate two variations of NORA as an ablation study. If we only consider term $T_{1}$ and term $T_{3}$ in the approximation (“NORA-$T_{1}$” in the tables), we only consider the task-specific influence of the embeddings but ignore the structural influence, which equals setting $k_{3}^{\prime}$ to zero in Equation 16. If we only consider term $T_{2}$ and term $T_{3}$ (“NORA-$T_{2}$” in the tables), we only consider the structural influence, which equals to setting $k_{3}^{\prime}$ to infinity in Equation 16. NORA outperforms both variants, demonstrating the benefit of ensembling the task-specific influence and structural influence. In some situations, only considering one of them could already have a good performance that is close to NORA. It indicates that sometimes the influence is dominated by the task-specific influence or the structural influence. Ensembling both of them provides us with the opportunity to appropriately utilize both of them.